# Navigating the Future of Federated Recommendation Systems with Foundation Models

### v1

@@ III-B2 Adapter
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adapter. This technique adds new lightweight layers while maintaining the original parameters of the pre-trained model, enabling fine-tuning and extension for specific tasks. This approach is suitable for multi-task learning and tasks performed in resource-limited environments. Adapters, such as LoRA [62] and QLoRA [63], have found wide applications in natural language processing [62] and computer vision tasks.
@@ IV Federated Recommendation Systems with Foundation Models
As described earlier in Fig. 3, a typical FRS typically consists of three stages: client model update, communication, and global aggregation. Integration with FM should also occur at these three stages.
@@ IV-A Client Model Update
In FRS, the clients have the following characteristics:
  • As previously described, the data needed for model training in FRS, such as user information and interaction history, are privacy-sensitive and are therefore required to stay on the user’s device, creating data silos.
  • Moreover, each user’s data is minimal compared to the total dataset, and each user only accesses a small portion of the item set, leading to data sparsity.
  • The data on the client is influenced by user preferences, such as user behavior and product preferences. The resulting data distribution often does not meet the independent and identically distributed (IID) assumption, leading to the challenge of data heterogeneity.
@@
  • Additionally, user devices are generally consumer products like mobile phones and personal computers, characterized by unstable communication and limited computational resources. This requires that the computational load of the models deployed on the client and the amount of information exchanged with the server be kept as low as possible.
FM is typically pre-trained on large, diverse datasets, acquiring a broad range of features and knowledge. This pre-training endows them with prior knowledge that allows for rapid adaptation to specific client data through fine-tuning. Consequently, these models can learn general feature representations, providing a certain level of adaptability to different data distributions. Therefore, clients can effectively fine-tune the model with a minimal amount of local data, achieving good performance on specific downstream tasks. Additionally, by fine-tuning foundation models locally, sensitive data does not need to leave the device, thereby enhancing data privacy.
@@
However, it is important to note that when FM is applied to data significantly different from the training distribution, performance degradation may occur. This issue, known as out-of-distribution generalization, represents a challenging aspect that FMs need to overcome. Moreover, if biases exist in the training dataset, FMs might learn and amplify these biases, leading to unfair outcomes across different data distributions. Although fine-tuning FMs requires significantly fewer resources than training from scratch, it usually still demands substantial computational resources for effective fine-tuning and updating. This requirement could limit their application on resource-constrained clients, particularly in FRS, where each user represents a client, thus potentially restricting the deployment of FM due to limited computational resources.
@@ IV-B Communication
Communicate Efficiency. Similar to FL, FFRS involves significant data transfer between numerous clients and a central server, making communication efficiency a critical factor in the duration of the entire learning process. To enhance communication efficiency, model compression techniques such as quantization or sparsification of model parameters can be utilized to reduce the volume of data transmitted. Periodic averaging is another strategy that reduces communication overhead by decreasing the frequency of model parameter uploads.
FM, due to its parameter sharing and hierarchical features, allow clients to transmit only fine-tuned parameter updates rather than the entire model’s parameters. This approach can significantly reduce communication overhead while maintaining or even enhancing the model’s performance on specific tasks. Similarly, since foundation models have already learned rich feature representations during the pre-training phase, they can adapt more quickly to new tasks. This means that in FRS, clients may achieve satisfactory performance with fewer iterations, thus reducing the frequency of communication.
Client Selection. In each iteration of FRS, considering resource limitations and privacy concerns, the central server may choose to communicate only with a subset of clients. Research indicates that clients should not participate in consecutive training rounds as attackers could potentially infer information about the client from transmitted gradients or model parameters. Client selection strategy determines which clients will participate in the current round based on various factors such as computational power, network stability, data diversity, and quality. Client selection can also be conducted through random sampling or based on statistical characteristics of the data to ensure the model learns from diverse data sources, enhancing its generalization ability.
Since FM already possesses a broad knowledge base and generalization capabilities, when selecting clients for model training, greater emphasis can be placed on data compatibility and coverage. Specifically, priority can be given to clients that can provide data types less represented or missing in the pre-training phase, thereby supplementing and enhancing the model’s performance in these areas. For instance, clients whose data can significantly improve the model’s performance on specific tasks can be considered more valuable. Incentive mechanisms can be set up to encourage the participation of these high-value data providers in the training process. However, it is important to ensure that the same client does not participate in consecutive training rounds to minimize the risk of privacy breaches.
Privacy Protection. A major advantage of federated learning is the ability to train models without compromising user privacy, making privacy protection a primary consideration in designing communication protocols. Privacy risks are mitigated by transmitting model updates, such as gradients or parameters, instead of raw data from clients to the server. Additionally, technologies like homomorphic encryption, secure multi-party computation (SMC), or differential privacy can further enhance privacy protection during communication.
@@ IV-C Global Aggregation
Global aggregation in FRS integrates model parameters independently trained by various clients to form a unified global model. This process not only enhances the accuracy and generalization of the recommendation system but also strengthens the system’s robustness by protecting client data privacy, optimizing resource usage, ensuring model synchronization and fairness, thereby improving recommendation performance and efficient resource utilization without compromising user privacy.
IV-C1 Typical Average
The Typical Average scheme, proposed with FedAvg [3], is the most common aggregation strategy where the server computes a simple arithmetic average of the model updates received from client devices. It’s popular due to its simplicity and effectiveness in many scenarios
IV-C2 Weighted Average
In the Weighted Average strategy, model updates are weighted based on certain criteria such as the volume of data on each client or the reliability of the data source, allowing more significant contributions to have a proportionally greater impact on the global model. This approach can be more effective than the typical average scheme, especially in non-IID data environments, where data distribution varies significantly across devices
IV-C3 New Future Methods
When integrating FM, exploring new aggregation technologies becomes particularly important to address the challenges posed by the scale and complexity of the models. Beyond the traditional weighted average method, some emerging aggregation strategies have shown potential in managing these challenges. For instance, the technique called Model Soups enhances model accuracy and robustness by averaging the weights of models that have been fine-tuned with different hyperparameters. Additionally, aggregation strategies based on the Mixture of Experts (MoE) utilize multiple specialized sub-models, each optimized for specific tasks or data types. These strategies dynamically adjust model weights based on data-driven methods, allowing for flexible adjustment of the aggregation process in response to real-time data changes, thus improving the overall model performance and adaptability. These emerging technologies not only improve the effectiveness of aggregation but also better address the challenges from large-scale distributed data, making them valuable aggregation strategies worth further research and application in the field.
Additionally, leveraging machine learning techniques to dynamically adjust aggregation methods based on real-time data quality and integrity assessments could improve both the robustness and efficiency of FRS.
@@ V Challenges
The integration of generative models into federated recommendation systems heralds a new frontier in personalized content delivery, promising to enhance user experiences with tailored suggestions while respecting privacy concerns. However, this promising union is not without its complexities. As we venture further into this domain, a myriad of challenges emerges, each posing unique hurdles that must be carefully navigated. These challenges span the spectrum from data heterogeneity and privacy concerns to communication overheads and resource-intensive computations. Understanding and addressing these issues is paramount to the successful deployment and operation of federated recommendation systems that are both effective and secure. In this section, we delve into each of these challenges, exploring their implications and discussing potential strategies for overcoming them.
V-A Data heterogeneity
The diversity in user-generated data can complicate the generative model’s ability to identify a common representation that accurately reflects all users’ preferences. This challenge necessitates the development of robust mechanisms capable of handling varied data inputs and providing nuanced, precise recommendations. Techniques such as transfer learning and multi-task learning can be employed to better adapt to the heterogeneity of data.
V-B Privacy and security risks
The collection of user data is essential for refining recommendation algorithms, but it also poses significant privacy and security risks. While federated learning offers a solution by processing data locally, generative models must be carefully designed to prevent the leakage of sensitive information through their generated outputs. Privacy-preserving methods like differential privacy and secure multi-party computation can be integrated to enhance data security.
V-B1 Memeber Inference Attacks
Membership inference attacks aim to determine whether specific data samples were used in training a model. Such attacks can reveal the presence of sensitive data, for example, by disclosing whether a patient’s medical records were used to train a disease prediction model. Under the FL setting, attackers might use model updates obtained from various participants to infer which data were used for training. Particularly in FRS with FM settings, the complexity and depth of the models make membership inference attacks more covert, thereby complicating defense efforts.
V-B2 Data Reconstruction Attacks
Data reconstruction attacks aim to rebuild or approximate the actual data used in training. This is typically achieved through optimization techniques such as model inversion and gradient matching, where attackers attempt to generate data samples similar to the training data by accessing the model’s outputs or gradients. Under the FL setting, although the original data does not leave the owner’s device, aggregated updates may still leak sufficient information to enable attackers to reconstruct the original data. Particularly when large-scale foundation models are used, this risk may be exacerbated due to the models’ high capacity for data representation.
V-B3 Poisoning Attacks
In FL, poisoning attacks represent a security threat aimed at disrupting or manipulating the learning process and outcomes through malicious modifications to data or model parameters. These attacks primarily take two forms: data poisoning attacks and model poisoning attacks. Specifically:
  • Untargeted Attacks: Untargeted attacks aim to disrupt the entire training process, causing the global model to fail to converge or significantly degrade in performance. These attacks are typically executed by injecting noisy data or incorrect information into the model training process. In a FL environment, attackers might submit model updates containing erroneous gradients or parameter updates, thereby disturbing the learning process of the aggregated global model. The challenge of these attacks lies in the ability of the attackers to pass the poisoned data through the FL system’s normal aggregation process without detection.
  • Targeted Attacks: In contrast, targeted attacks aim to cause the model to produce incorrect outputs for specific inputs, without affecting the model’s performance on most other inputs. This type of attack is usually carried out by injecting a small amount of carefully designed poisoned samples into the training data. These samples contain specific triggers that, when encountered by the model, lead to predetermined incorrect outputs. This attack method is particularly covert in FL, as attackers can embed these triggers during local training, and these modifications may only affect a small part of the model, making them difficult to detect in the global model.
V-C Communication Cost
Federated systems necessitate ongoing communication between users and servers to synchronize model updates. The large parameter size of generative models can significantly increase the demand for communication resources. To address this, efficient communication protocols and model compression techniques can be utilized to reduce the bandwidth required for updates.
V-D Data Scarcity
The scarcity of user data is a significant challenge, particularly for new or less active users, as it can hinder the model’s ability to learn accurate preferences. To combat this, data augmentation strategies and synthetic data generation can be employed to enrich the training data and improve model generalization.
V-E High Resource Demand
Generative models are computationally demanding, requiring substantial storage and processing capabilities. This can be particularly challenging in resource-constrained environments, such as mobile devices. Optimizing model architectures and leveraging distributed computing can help alleviate the strain on resources.
V-F Problematic Synthetic Data Generated
The quality of synthetic data generated by generative models is crucial. Low-quality synthetic data can negatively impact the model’s performance if used in training. Ensuring the synthetic data closely resembles real data distributions is essential for maintaining the integrity of the recommendation system.
V-G Robustness Consideration
Generative models within federated recommendation systems must be robust against data anomalies and capable of maintaining stable performance even when faced with shifts in data distribution. This requires the implementation of robust learning algorithms that can adapt to changes in the data landscape without significant degradation in performance.

### v2

@@ III-B2 Adaptation
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adaptation. This technique adds new lightweight layers while maintaining the original parameters of the pre-trained model, enabling fine-tuning and extension for specific tasks [83, 41]. This approach is suitable for multi-task learning and tasks performed in resource-limited environments, such as few-shot learning [84]. Adaptations have found wide applications in natural language processing and computer vision tasks [85, 86, 87] . According to [84], adaptation methods are categorized into three main types:
  • Prompt-based Fine-tuning Adaptation Methods are inspired by the principles of prompt learning in NLP, adapts the model to a specific downstream task by using a portion of the fixed text input as a learnable vector and fine-tuning it using downstream task data. For instance, CoOp [88] introduces the concept of prompt learning into the adaptation of downstream tasks for multimodal pre-trained foundation models by automatically constructing contextual prompts using learnable word embeddings. CoCoOp [89] builds upon this by creating a meta-network that learns image features, which are then combined with prompt vectors to enhance the generalization performance of CoOp on new category data.
  • Adapter-based Fine-tuning Adaptation Methods entail the incorporation of small, trainable modules called Adapters into a pre-trained model. These adapters are subjected to fine-tuning, rather than the entire model, thereby enabling the model to adapt to novel tasks with minimal alterations to its original parameters. CLIP-Adapter [87] adds simple adapter structures within the CLIP model that are fine-tuned using a small dataset. This method allows the model to adapt to new tasks by learning task-specific features without the need for extensive retraining of the entire model. CALIP [90] enhances the cross-modal interaction capabilities of adapters by integrating text and image features using attention maps and incorporating two trainable linear layers before and after this integration. LoRA [85] adjusts the weights of a pre-trained model by multiplying the weight matrices with a low-rank matrix, allowing for fine-tuning of the model without significantly increasing the number of parameters.
  • External Knowledge-based Adaptation Methods integrate external knowledge (e.g., from knowledge graphs or additional language model queries) into the adaptation process. This is particularly advantageous for enhancing the model’s comprehension and efficacy in tasks where domain-specific knowledge is vital but data is scarce. CuPL [91]enhances zero-shot classification performance by asking questions to a LLM and generating multiple descriptive statements for each category, enriching the semantics of the categories. LaBo [92] uses GPT-3 to generate a large pool of candidate feature descriptors. It then selects the optimal subset of descriptors for classification decisions, considering the distinctiveness of features for other categories and the coverage for the current category.
@@ IV Federated Recommendation Systems with Foundation Models
As described earlier in Fig. 2, a typical FRS typically consists of three stages [27]: client model update, communication, and global aggregation. Integration with FM should also occur at these three stages.
@@ IV-A Client Model Update
In FRS, clients have the following characteristics [26, 27]:
  • As mentioned earlier, a major advantage of FRS is privacy. The data required for model training in FRS, such as user information and interaction history, is privacy sensitive and therefore must remain on the user’s device, creating data silos. This approach not only complies with data privacy regulations (such as GDPR), but also increases user trust. However, this means that each user has a relatively small amount of data compared to the entire dataset, and each user only accesses a portion of the item set, leading to data sparsity.
  • Each client independently trains the model on its local data. This allows the model to better adapt to each user’s unique preferences and behaviors, improving personalization and accuracy of recommendations. The resulting data distribution often does not satisfy the independent and identically distributed (IID) assumption, leading to the problem of data heterogeneity.
@@
  • In addition, user devices are typically consumer products such as mobile phones and personal computers, which are characterized by unstable communication and limited computational resources. This requires keeping the computational load of the models deployed on the client and the amount of information exchanged with the server as low as possible.
FM is typically pre-trained on large, diverse datasets, providing it with rich knowledge and strong representational capabilities. This pre-training provides it with prior knowledge that allows it to quickly adapt to specific customer data through fine-tuning [18]. As a result, FM can learn general feature representations that provide a degree of adaptability to different data distributions. Applying these models during the client update phase in FRS can significantly improve the performance of local models. Specifically:
  • FM can apply knowledge learned from large corpora to local data through transfer learning. For example, a pre-trained FM can be effectively fine-tuned on a small amount of local data in the client to adapt to specific user behavior data, thereby achieving good performance in specific downstream tasks such as providing more accurate recommendations [38]. In addition, by fine-tuning the base model locally, sensitive data does not need to leave the device, protecting user privacy [93]. As a result, customers can achieve better model performance with limited data.
  • FM can also capture complex user preferences and behaviors. For example, FM has strong semantic understanding capabilities that allow them to better understand users’ search queries [63, 64], comments [37], and other textual data [94] to provide recommendations that better meet user needs. In addition, as user behavior changes, FM can quickly adapt and adjust recommendation strategies to ensure that recommendation results remain consistent with the user’s current interests and needs. By applying FM during the client update phase, a higher level of personalized recommendation can be achieved.
@@
However, it is important to note that when FM is applied to data that differs significantly from the training distribution, performance degradation can occur. This problem, known as out-of-distribution generalization [93], is a challenge that FM must overcome. In addition, if there are biases in the training data set, FM may learn and reinforce these biases, leading to inequitable results across different data distributions [95, 96, 97]. Although fine-tuning FM requires significantly less resources than training from scratch, it still typically requires significant computational resources for effective fine-tuning and updating [18, 36]. This requirement may limit its use on resource-constrained clients, especially in FRS where each user represents a client, potentially limiting the use of FM due to limited computational resources.
@@ IV-B Communication
In FRS, communication between clients and the server is a critical phase. The primary task at this stage is to transmit model update information from each client to the central server while ensuring data security and communication efficiency during the transmission process. In traditional FRS frameworks, communication between clients and the server typically involves transmitting large amounts of update information, which may include model parameters, gradients, or other relevant statistical data. To ensure efficient and secure communication, common strategies include data compression [98, 99, 100], encrypted transmission [101, 102], and secure aggregation [103, 104, 105]. However, with the increasing data scale and higher privacy protection requirements, these traditional methods face significant challenges [36].
FMs, especially LLMs, possess strong representation capabilities, enabling them to extract key features from model updates through efficient data representation [55, 106] and semantic compression techniques [107, 108]. By removing redundant information and unnecessary data, and retaining only essential information, FM can transform this into a more compact representation, significantly reducing the data transmission volume between clients and the server. Optimizing communication protocols [109] and enhancing data compression and security during this process can markedly improve the overall performance of the system.
@@ IV-C Global Aggregation
In FRS, the global aggregation phase of the server is responsible for combining model updates from all clients into a single global model. This phase is of critical importance in ensuring the overall performance and effectiveness of personalized recommendations. Server global aggregation is characterized by the following key features [110]: First, server global aggregation must handle updates from a large number of clients that contain different data and user behaviors. To ensure the overall quality of the model and the accuracy of the recommendations, the server must effectively integrate this diverse data. Second, the global aggregation process must be sufficiently robust to handle potentially noisy data and anomalous updates. Finally, the scalability of the server’s global aggregation is critical, especially when dealing with a large number of clients and massive amounts of data, requiring the system to efficiently process and aggregate updates. In each iteration, given resource constraints and privacy concerns, the central server may only communicate with a subset of clients. Researches [111, 112] indicate that clients should not participate in successive training rounds, as attackers could potentially infer information about the client from transmitted gradients or model parameters. The client selection strategy determines which clients will participate in the current training round based on factors such as computational power, network stability, data diversity, and data quality [113]. For example, client selection can also be done by random sampling or based on statistical characteristics of the data to ensure that the model learns from diverse data sources, thereby increasing its generalization ability.
Context-aware aggregation represents a critical application of FM in global server aggregation. Traditional aggregation methods, such as naive weighted averaging proposed in FedAvg [3], may not fully exploit the contextual information present in the data. FM, however, is capable of intelligently aggregating updates by understanding the semantics and contextual relationships of each participant’s updates [70]. To illustrate, when FM receives customer updates, it analyzes the contextual information of each update, such as user behavior patterns, preferences, and history. This information enables FM to better comprehend the significance and value of each update and to assign distinct weights accordingly [32]. This context-aware aggregation method can more accurately reflect user preferences and behaviors, thereby enhancing the overall performance of the model. Dynamic weighting permits the system to adapt more effectively to users’ personal needs and to provide recommendations that are more aligned with their interests.
With its extensive knowledge base [114], FM can utilize this knowledge to make more informed decisions during the aggregation process. Knowledge-based aggregation strategies can capitalize on FM’s knowledge and reasoning capabilities to enhance the efficacy of aggregation. To illustrate, in the event that certain customers lack sufficient data, FM can leverage pertinent information from its knowledge base to infer relationships and patterns between data, thereby supplementing and improving aggregation. The key concept of Mixture of Experts (MoE) [115] is that different expert models specialize in different parts of a problem. By appropriately weighting and combining their outputs, better overall performance can be achieved. The MoE architecture has a similar function to aggregation strategies in FRS. Therefore, applying this architecture to the aggregation strategy in FRS may result in improved recommendation performance.
In terms of anomaly detection, FM also performs exceptionally well [116, 117]. In FRS, client updates may be influenced by various factors, leading to noisy or abnormal data [29]. If these abnormal updates are directly used in global aggregation, they may affect the overall quality of the model. FM can utilize its strong anomaly detection capabilities to analyze client updates, identify and handle these anomalies, ensuring the robustness of global aggregation. For instance, if a client’s data update exhibits a notable divergence from the majority of client updates, it may be regarded as noise or an anomaly. FM has the capacity to elect to disregard or assign a diminished weight to it, thus reducing its deleterious impact on the global model and ensuring the stability and robustness of the global model. Liu et al. [37] observed in their study that FM is capable of effectively filtering out noisy data by identifying aberrant patterns, thereby enhancing the stability and reliability of the aggregation process.
This application direction has considerable potential for future development and is worthy of further research and investigation.
@@ V Challenges
The application of FM to FRS marks the advent of a new frontier in personalized content recommendation. This approach promises to further enhance user experience by employing more powerful models than traditional deep models while protecting user privacy. However, this promising combination comes with many challenges, involving various aspects such as data privacy and security, data sparsity and sample imbalance, model training and optimization, model interpretability and transparency, and so on. It is therefore of the utmost importance to gain a comprehensive understanding of these issues and to develop effective strategies to address them. This is essential if FRS is to be successfully deployed and operated in a secure yet powerful manner. In this section, we will examine these challenges in detail, explore their implications and discuss potential strategies to overcome them.
V-A Data-Related Challenges
  • Data Privacy: In a federated setting, the protection of user privacy is of paramount importance. The model must ensure that training and recommendations are performed without disclosing personal data. However, FMs, such as GPT-3, have the potential to memorize and reproduce data from the training set, thereby potentially leaking sensitive information. Additionally, if the generated data is too similar to the original data, there is a risk of privacy infringement. This may involve the use of additional encryption techniques to safeguard user data. Differential privacy [118] protects individual data by adding noise during model training, thus reducing the risk of data leakage while maintaining model performance. Homomorphic encryption [119] allows computations on encrypted data, enabling model training without decrypting the data, effectively preventing data theft during transmission and processing. Furthermore, machine unlearning techniques [120] can remove specific user data from the foundational model, ensuring that these data no longer influence the model’s predictions and outputs. This approach is particularly effective in the context of user deletion requests or in the context of compliance with privacy regulations such as the GDPR.
  • Data Security: The integration of FRS and FM also necessitates the equal consideration of data security. Participants may be subjected to malicious attacks, including Member Inference Attacks [121], Data Reconstruction Attacks [122], and Poisoning Attacks [123], which could compromise the model or alter the data. Consequently, ensuring the integrity and confidentiality of data during transmission and storage represents a significant challenge. The application of secure multi-party computation (MPC) [124] can distribute computational tasks among multiple participants, ensuring that no single party can obtain complete data, thereby enhancing security. Additionally, blockchain [125, 126] can be used to record data access and operation logs, ensuring data integrity and traceability. This also helps improve the transparency and security of the system.
  • Data Sparsity and Scarsity: RS frequently encounters the challenge of data sparsity, where there is limited interaction data between users and items. This leads to decreased model performance. This issue is particularly pronounced in a federated setting, where each participant has limited data. In specific domains or tasks, insufficient data can cause severe data scarcity, further restricting model training and performance. Although FM has the ability to perform zero-shot and few-shot learning, they still need optimization to handle extremely sparse data. It is therefore of great importance to utilise data augmentation techniques [127, 128] in order to generate a greater quantity and variety of training samples, thereby increasing the overall data set. Alternatively, knowledge transfer [129] between different domains can be employed in order to leverage the rich data sets from other areas, thus alleviating the issue of data sparsity and improving the overall model performance in the current domain. Despite the ability of generative FM [130, 131] to produce a vast quantity of synthetic data, the quality of this data may not be guaranteed. If the FM utilises low-quality or biased data sources, the generated data may inherit these issues and even exacerbate existing biases and errors [132]. Wang et al. [22] found that generated data is not always beneficial for representation learning. For example, in contrastive learning, the optimal performance was achieved when the mix ratio of real to generated data was approximately 10:1. Consequently, when generating synthetic data, it is of paramount importance to establish quality standards and utilise evaluation metrics to assess data quality. Furthermore, it is essential to implement quality control mechanisms to filter out low-quality or biased data.
  • Sample Imbalance: In a federated setting, there are significant differences in data size and distribution among different clients [133, 134, 135]. Similarly, in RS, item labels and user behaviors often exhibit a long-tail distribution [136]. These issues are referred to as sample imbalance, where part of classes have much more samples than others in the dataset. This can cause the model to overfit the majority classes while having poor generalization for minority classes, thus affecting the model’s training effectiveness. To address this, additional methods are required to balance these differences and ensure the model’s fairness and accuracy. Resampling techniques typically involve either oversampling or undersampling to balance the dataset [137]. Oversampling increases the number of minority class samples, while undersampling reduces the number of majority class samples. Weighted loss functions can assign higher weights to minority class samples during model training, enhancing the model’s ability to recognize minority classes. This approach can enhance the fairness and accuracy of the model without modifying the data distribution.
V-B Model-Related Challenges
  • Model Synchronization: In a federated setting, models from various participants frequently communicate with the server for synchronization, resulting in high communication costs and complex system management issues. This is particularly problematic for large FM due to their vast number of parameters [18], which require significant computational resources and time for transmission and synchronization. The use of gradient compression techniques [138] to reduce the amount of data transmitted during model synchronization can effectively lower communication costs and enhance synchronization efficiency. Furthermore, the implementation of asynchronous update strategies [139], which permit participants to independently update their models and periodically perform global synchronization, can reduce waiting times during synchronization and enhance the overall efficiency of the system.
  • Model Heterogenity: Model heterogeneity [140] is typically attributable to a multitude of factors, including disparate types, sizes, and architectures of models among various clients. This can be attributed to differences in client hardware, data distribution, or business needs. In certain instances, different clients may utilize distinct model structures. For instance, some clients may employ simple linear models, while others may utilize complex deep learning models [141]. This renders it challenging to deploy or fine-tune a unified foundation model in a federated setting. The design of adaptive training algorithms [142, 143] enables the dynamic adjustment of model parameters based on the computational capabilities and data characteristics of each participant, thereby enhancing the model’s adaptability and performance in heterogeneous environments. Furthermore, knowledge transfer [47] among different participants facilitates the sharing of insights and the mitigation of data and computational resource disparities, resulting in an improvement in the overall performance of the model.
  • Model Interpretability: FM is often represented by black-box models, which makes it challenging to comprehend their internal workings. This lack of interpretability can give rise to issues of trust and compliance in practical applications. In FRS, model interpretability is of paramount importance for establishing user trust and meeting regulatory requirements [144]. Therefore, it is vital to employ explainable AI techniques [145] to provide transparent and understandable explanations for the decisions made by FM. For instance, attention mechanisms [146] and feature importance analysis [147] can be utilized to elucidate model decisions. Furthermore, the utilisation of generative FM [148] to generate natural language explanations for recommendations can provide users with the context and rationale behind the model’s decisions, thereby enhancing user trust and acceptance of the model.
  • Model Transparency: The concept of model transparency entails the open and auditable process of model decision-making. In the context of FRS, transparency plays a pivotal role in fostering user trust, necessitating the assurance of model fairness and non-discrimination. This encompasses the monitoring of model training processes, the validation of model updates, and the tracing of generated data. A lack of transparency can diminish the credibility of the model and increase compliance risks. FRS with FM should have transparent training protocols [149] that record and monitor each step of the model training process. This ensures transparency, thereby enhancing system credibility and compliance. Similarly, establishing a detailed logging system [48] to record every model update and operation ensures traceability, helping to identify and correct issues within the model and improve overall system transparency.